OpenStudy (anonymous):

Given that three angles of a triangle are represented by x, x+5, and 3x, what are the measures of those angles if the sum of three angles of a triangle are always equal to 180 degrees

OpenStudy (anonymous):

Since the sum of them = 180, then add them all up : x + 3x + x+5 = 180 5x = 175 x = 35 Substitute x in all you'll get : 1) x = 35 degrees 2) x + 5 = 35 + 5 = 40 degrees 3) 3x = 3(35) = 105 degrees ^_^

OpenStudy (anonymous):

Thank you.. I asked another question!

OpenStudy (anonymous):

most welcome ^_^ I have answered it before.